OpenStudy (anonymous):

A(-1,3), B(2,5) Draw vectors AB and BA and give vectors in component form equivalent to each of these vectors.

OpenStudy (unklerhaukus):

\[\vec{\boldsymbol v} =(v_x,v_y)=v_x\,\hat{\mathbf x}+v_y\,\hat{\mathbf y}\]

OpenStudy (mathmale):

Given these two points, there is only one vector from A to B. My take is that you are to provide the horiz. and vertical components of the vector from A to B.

OpenStudy (mathmale):

Might be worthwhile to double check to ensure you have copied (typed in here) the problem and its instructions exactly as you found them.

OpenStudy (anonymous):

I don't think the question was to find the vector from A to B

OpenStudy (anonymous):

\[\left(\begin{matrix}-1 \\ 3\end{matrix}\right)\] \[\left(\begin{matrix}2 \\ 5\end{matrix}\right)\]